Ok I have tried to figure out what they want me to do on my own but I keep getting a wrong answer. The problem is: If Lynn can type a page in P minutes, what piece of the page can she do in 5 minutes?


If you can just tell me what they are asking me to do(multiply, divide, etc..) Then I can figure it out.


Thank you, Klampo

Found 2 solutions by mananth, dnanos:
Answer by mananth(16949)

Ok I have tried to figure out what they want me to do on my own but I keep getting a wrong answer. The problem is: If Lynn can type a page in P minutes, what piece of the page can she do in 5 minutes?
p minutes -------------- 1 page
5 minutes ----------------x
Direct variation
..
p/5 = 1/x
x = 5/p of the page
..
m.ananth@hotmail.ca


Answer by dnanos(83)

The whole page in P minutes, and the piece of it in 5 minutes.
You must make a fraction 5/P , (that is you divide 5 by P)
and this is the solution.